At a distance H above the surface of a planet, the true weight of a remote probe is eight percent less than its true weight on the surface. The radius of the planet is R. Find the ratio H/R.
H
R
=
1.24
1
Expert's answer
2019-10-16T09:38:45-0400

Due to the Newton's law of gravity the gravitational acceleration is (we assume "r \ge R" , where "R" - radius of the planet, also we assume that the planet has a spherical form)


"\left| {\vec g} \right| = {{GM} \over {{r^2}}}"

where "G" - gravitational constant, "M" - mass of the planet and "r" - distance. Then at the surface of the planet


"g = {{GM} \over {{R^2}}}"

and at the distance H above the surface


"g' = {{GM} \over {{{(R + H)}^2}}}"

According that the weight is "\vec P = m\vec g" and the mass of the body "m" does not change, we know from the conditions of the problem that "g' = kg = 0.92g" .

Then we have


"{{GM} \over {{{(R + H)}^2}}} = k{{GM} \over {{R^2}}}""{(R + H)^2} = {1 \over k}{R^2}""{(1 + {H \over R})^2} = {1 \over k}"

and (we reject 'unphysical' root)

"{H \over R} = {1 \over {\sqrt k }} - 1"

In our case we get


"{H \over R} = {1 \over {\sqrt {0.92} }} - 1 \approx 0.0426"